Circular No. 6613

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bp IN NGC 4680 P. Marples, Loganholme, Queensland, reports the visual discovery with

a 0.31-m telescope by R. Evans (Hazelbrook, N.S.W.) on Apr. 6.52 UT of a supernova (mag ∼ 13.8) located 10′′ west and 40′′ south of the center of NGC 4680 (α = 12^{h}46^{m}55^{s}, δ = −11^{o}38′.3, equinox 2000.0). No star appears in this position on the Palomar Sky Survey. Marples confirmed the existence of a new star visually on Apr. 6.6.

M. Phillips, Cerro Tololo Interamerican Observatory (CTIO), obtained a spectrogram of SN 1997bp on Apr. 7.1 UT with the CTIO 1.5-m telescope; the spectrum shows this to be a type-Ia supernova caught near maximum light. Offsets from the nucleus were measured at the telescope to be ∼ 13′′.0 west and 19′′.5 south.

D. J. Schlegel, University of Durham, obtained a spectrogram (range 473–724 nm) of SN 1997bp on Apr. 7 with the 4-m reflector at CTIO. Inspecting the uncalibrated data, A. V. Filippenko and D. C. Leonard, University of California at Berkeley, conclude that the object is a normal supernova of type Ia within 1 week of maximum brightness; the red Si II absorption trough is very deep.

L. Wang and J. C. Wheeler, University of Texas, write that a spectrogram (range 450–720 nm) obtained with the McDonald Observatory 2.7-m reflector (+ SPOL) on Apr. 7.2 UT also confirm that SN 1997bp is of type Ia at maximum light; the spectrum is very similar to the type-Ia SN 1981B at maximum light.

P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, communicate that spectra were obtained at the Multiple Mirror Telescope on Apr. 7.4 UT. Strong, broad Si II absorption is seen at 606.5 nm and Ca II at 368.6 nm. The spectrum is again confirmed to be that of a type Ia supernova at maximum. N. Grogin obtained CCD images of NGC 4680 with the Whipple Observatory’s 1.2-m telescope on Apr. 7.3, showing that SN 1997bp is located 14′′.6 west and 20′′.0 south of the galaxy center at magnitude V = 14.2.

SUPERNOVA 1997bn IN UGC 4329 W.-d. Li, Beijing Astronomical Observatory, reports a precise position

of SN 1997bn (cf. IAUC 6612), measured from unfiltered CCD frames taken on Apr. 4: α = 8^{h}19^{m}02^{s}.20, δ = +21^{o}11′01′′.7 (equinox 2000.0), which is 4′′.7 east and 7′′.2 south of the galaxy’s center.

1997 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green